# Continual Adversarial Defense

### 15 Dec 2023 v1

@@ Abstract
In response to the rapidly evolving nature of adversarial attacks on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that can generalize to all types of attacks, including unseen ones, is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks used by many attackers. The defense system needs to upgrade itself by utilizing few-shot defense feedback and efficient memory. Therefore, we propose the first continual adversarial defense (CAD) framework that adapts to any attacks in a dynamic scenario, where various attacks emerge stage by stage. In practice, CAD is modeled under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial images. We leverage cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Experiments conducted on CIFAR-10 and ImageNet-100 validate the effectiveness of our approach against multiple stages of 10 modern adversarial attacks and significant improvements over 10 baseline methods. In particular, CAD is capable of quickly adapting with minimal feedback and a low cost of defense failure, while maintaining good performance against old attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.  [^1] Code at https://github.com/cc13qq/CAD
Figure 1: In a dynamic environment, the target model $f_{t}$ faces a diverse range of attacks at different stages and is protected by the defender. During the $i$-th stage, the defender $f_{i}$ leverages few-shot defense feedback provided by the security department and users of the target model to adapt to new attacks. The defender also needs to efficiently use memory to store the cache, retaining knowledge of past attacks.
@@ 1 Introduction
From an application perspective, it is not realistic to design a defense method that can generalize to all types of attacks, including unseen ones. The truth is, the increasing number of defense strategies naturally stimulates the development of new attack algorithms [32]. As a result, the defense system encounters a dynamic environment characterized by numerous unique attacks with higher success rates in both white-box and black-box settings [16]. The target model is threatened by many attackers with various attacks, just like the defense scenario against Deepfake [10].
To simulate a dynamic environment more realistically, we propose a defense scenario that considers practicality. As illustrated in Fig. 1, the attacker employs diverse attacks at different stages to target the model, while the defender adapts to emerging attacks by leveraging few-shot defense feedback. In this paper, we use adversarial examples with ground truths as the defense feedback, which is provided by the security department and users of the target model in real-world situations. In this ongoing “arms race,” the attacker operates within the gray-box setting [32], where they possess knowledge of the classifier’s architecture and have access to training data but are unaware of the defense mechanism in place. On the other hand, the defender has the same training data and access to the target model, but also possesses full knowledge of the initial attack.
Taking real-world considerations into account, we propose four principles for the scenario: (1) Continual adaptation to new attacks without catastrophic forgetting. As defenders, it’s essential to adapt to various new attacks at different stages while retaining knowledge of previous ones. (2) Few-shot adaptation. Increased feedback indicates more successful attacks on the target model. Therefore, we cannot afford too many successful attacks before upgrading the defense. (3) Memory-efficient adaptation. Over time, a continuous influx of attacks results in accumulating defense feedback, potentially leading to memory constraints. In practical terms, we may not have sufficient memory capacity to withstand this. (4) High accuracy in classifying both clean and adversarial images. A robust defense should not compromise the interests of those it protects. Hence, maintaining high classification accuracy on both clean and adversarial images is crucial.
@@
$\bullet$ In order to simulate the rapidly evolving adversarial attacks against image classifiers, we introduce, for the first time, a dynamic scenario of various attacks that emerge in different stages. The defender is required to adapt to new attacks using few-shot defense feedback and efficient memory, while also retaining knowledge of previous attacks and upholding high performance on both clean and adversarial images.
@@
$\bullet$ Extensive experiments on CIFAR-10 and ImageNet-100 validate the effectiveness of CAD against multiple stages of 10 modern adversarial attacks using few-shot feedback and efficient memory, and demonstrate significant improvements over 10 baseline methods.
@@ 2.2 Adversarial Defense
Certainly, both adversarial training and purification methods have shown limitations in robustness, leading to reduced classification performance on clean images. Furthermore, the dynamic nature of the environment introduces a diverse array of unique attacks employed by multiple attackers. In this paper, we propose Continual Adversarial Defense (CAD) framework as a solution to effectively combat various attacks emerging stage by stage, while maintaining strong performance in classifying clean images.
@@ 2.3 Continual Learning
Continual learning (CL) [21, 28, 19, 18, 20, 25, 22, 24, 46, 23] aims to learn from a sequence of new classes without forgetting old ones and attracts much attention to various computer vision tasks. Many works have been proposed for CL: by introducing different knowledge distillation (KD) losses to consolidate previous knowledge [28], by preserving a small number of old class data (exemplars) or augmented data (Zhu et al., 2021) to recall the old class knowledge [49], and by expanding the network capacity for new class data or freezing partial network parameters to keep the old class knowledge [48]. In recent years, some methods have aimed to solve the CL problem without depending on preserving data (called none-exemplar) [51], and some methods have attempted to learn new classes in a few-shot scenario [43] in which only a small number of new class data is gainable. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense feedback and efficient memory for adaptation.
@@ 3 Threat Model and Defense Principles
The defense system operates within a dynamic environment characterized by a multitude of unique attacks since the growing number of defenses naturally encourages the invention of new attack algorithms. Once frustrated by an existing defense, attackers undoubtedly resort to deploying various attacks in pursuit of breakthroughs. In response to this reality, we simulate the dynamic environment and put forth a practical threat model. In this threat model, attackers deploy different attacks in various stages to assault the target model, and the defender must adapt to evolving attacks by utilizing the defense feedback, just like an ”arms race”. The defense feedback is typically provided by security services and users of the target model upon the discovery of successful attacks during screening or cleanup efforts in the real world, and plays a critical role in informing and shaping the defender’s response to emerging threats. In this paper, we use adversarial examples with ground truths as the defense feedback. We formulate this attack vs defense scenario as follows.
Our scenario is established on the $N$-way $K$-shot classification paradigm. The attacker attacks under the gray-box setting [32] in which the attacker knows the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage. At the $i$-th stage, where $i=1,2,\cdots,T$, the attacker knows the failure of previous attacks and turns to a new attack $A_{i}(\cdot)$. Meanwhile, the defender receives a set of feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
@@
Figure 2: Illustration of CAD. Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate overfitting brought by future fine-tuning. And then, a weight estimator model $f_{p}$ is trained for the model ensemble. In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense feedback. To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process. After all, the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator to maintain high accuracy of both clean and adversarial images.
@@
In the dynamic environment, it is not uncommon for certain attackers to employ the same attack across different stages. Therefore, the defense mechanism must be capable of adapting to a range of new attacks in each stage while also retaining knowledge of previous ones.
@@
Over time, the ongoing influx of attacks results in accumulating defense feedback, potentially leading to the memory constraints. In practical terms, the defender may not have sufficient memory space to accommodate this accumulation. Consequently, storing all received defense feedback for adaptation is impractical and should be avoided.
@@
The previous defense strategies have led to a sacrifice in the classification accuracy of clean images, resulting in performance degradation. This decline could potentially have a negative impact on crucial real-world business operations. As a result, it is imperative for a defender to prioritize maintaining performance on both clean and adversarial images.
@@ 4.1 Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attack using defense feedback $\mathcal{A}_{\mathrm{train}}^{i}$. Inspired by continual-learning (CL) [28], we regard classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
Following [21], we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ at the $i$-th stage. The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$ and $N_{i}=N\times(i+1)$ is number of classes in stage $i$ after expansion. Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^2] We omit the networks’ parameter in formulas:
@@ 4.2 Embedding Reservation for Few-Shot Adaptation
First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as ‘virtual classes [43], where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
@@ 4.3 Prototype Augmentation for Memory Efficiency
In response to Principle 3, we use prototype augmentation to efficiently use memory.
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [50]. Many CIL methods store a fraction of old data to jointly train the model with current data to address this issue [28]. However, preserving old feedback may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation [51] to maintain the decision boundary of previous stages, without saving any feedback. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:

### 13 Mar 2024 v2

@@ Abstract
In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on. The defense system must gather online few-shot defense feedback to promptly enhance itself, leveraging efficient memory utilization. Therefore, we propose the first continual adversarial defense (CAD) framework that adapts to any attacks in a dynamic scenario, where various attacks emerge stage by stage. In practice, CAD is modeled under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial images. We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Experiments conducted on CIFAR-10 and ImageNet-100 validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods. In particular, CAD is capable of quickly adapting with minimal feedback and a low cost of defense failure, while maintaining good performance against previous attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.
Keywords:
Adversarial Defense Continual Learning
Figure 1: In a dynamic environment, the target model $f_{t}$ encounters a wide spectrum of attacks across various stages while being safeguarded by the defender. During the $i$-th stage, the defender leverages few-shot defense feedback, which is collected by the online defense system or provided by the security department and users of the target model, to adapt to emerging threats. Furthermore, the defender is required to efficiently manage memory utilization to store the cache, preserving insights gained from past attacks.
@@ 1 Introduction
From an application perspective, it is not realistic to design a defense method that can generalize to all types of attacks, including unseen ones. The truth is that the increasing number of defense strategies naturally stimulates the development of new attack algorithms [32]. As a result, the defense system encounters a dynamic environment characterized by numerous unique attacks with higher success rates in both white-box and black-box settings [20]. The target model is threatened by many attackers with various attacks, just like the defense scenario against Deepfake [13].
Inspired by online learning, the crucial approach to addressing dynamic scenarios lies in embracing a perpetually evolving defense strategy. Consequently, deploying a defense system on the cloud that continuously gathers attack examples from newly emerged adversarial algorithms and promptly adapts to these attacks represents a promising method for responding to such scenarios. As depicted in Fig. 1, a variety of attacks are detected across different stages, with the defender adapting to emerging threats by utilizing few-shot defense feedback. In this paper, we employ adversarial examples with ground truths as the defense feedback, sourced either from the online defense system’s crawling efforts or provided by the security department and users of the target model in real-world scenarios. The defender undergoes training with an initial attack and shares the same training data and architecture as the target model. Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
Taking real-world considerations into account, we propose four principles for the scenario: (1) Continual adaptation to new attacks without catastrophic forgetting. As defenders, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters. (2) Few-shot adaptation. The escalation in feedback signifies a greater potential for attacks on the target model. Therefore, we cannot tolerate an excessive number of attacks before adaptation. (3) Memory-efficient adaptation. Over time, a continuous influx of attacks leads to the accumulation of defense feedback, potentially posing memory constraints. In practical terms, we may not have sufficient memory capacity to withstand this pressure. (4) High accuracy in classifying both clean and adversarial images. A robust defense should not compromise the interests of those it protects. Hence, ensuring high classification accuracy on both clean and adversarial images is paramount.
@@
$\bullet$ To withstand the rapidly evolving adversarial attacks targeting image classifiers, we introduce, for the first time, a dynamic scenario featuring diverse attacks emerging at different stages to continuously enhance the defense. The defender must adeptly adapt to novel attacks using few-shot defense feedback and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial images.
@@
$\bullet$ Extensive experiments on CIFAR-10 and ImageNet-100 validate the effectiveness of CAD against multiple stages of 10 modern adversarial attacks using few-shot feedback and efficient memory and demonstrate significant improvements over 10 baseline methods.
@@ 2.2 Adversarial Defense
Indeed, both adversarial training and purification methods have exhibited limitations in robustness, resulting in diminished classification performance on clean images. Moreover, the dynamic environment introduces a multitude of unique attacks, rendering a single defense inadequate. In this paper, we present the Continual Adversarial Defense (CAD) framework as a solution to effectively counter various attacks emerging stage by stage, while ensuring robust performance in classifying clean images.
@@ 2.3 Continual Learning
Continual learning (CL) [22, 27] aims to learn from a sequence of new classes without forgetting old ones and attracts much attention to various computer vision tasks. Many works have been proposed for CL: by introducing different knowledge distillation (KD) losses to consolidate previous knowledge [27], by preserving a small number of old class data (exemplars) or augmented data (Zhu et al., 2021) to recall the old class knowledge [49], and by expanding the network capacity for new class data or freezing partial network parameters to keep the old class knowledge [47]. In recent years, some methods have aimed to solve the CL problem without depending on preserving data (called none-exemplar) [51], and some methods have attempted to learn new classes in a few-shot scenario [43] in which only a small number of new class data is gainable. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense feedback and efficient memory for adaptation.
@@ 3 Threat Model and Defense Principles
The defense system is deployed online and functions within a dynamic environment marked by a plethora of unique attacks, driven by the proliferation of defenses that naturally prompt the development of new attack algorithms. As time progresses, numerous new attacks, accompanied by attack examples, will emerge on the internet, aiming to exploit vulnerabilities in the defense system and achieve breakthroughs. In response to this reality, we simulate the dynamic environment and propose a practical threat model. In this model, distinct attacks are identified across various stages, requiring the defender to adapt to evolving threats through the utilization of defense feedback. Typically, defense feedback is obtained through crawling efforts of the online defense system or is provided by security services and users of the target model. This feedback is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world. In this paper, we use adversarial examples with ground truths as the defense feedback. We formulate this attack vs defense scenario as follows.
@@
Figure 2: Illustration of CAD. Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate overfitting brought by future fine-tuning. Then, a weight estimator model $f_{p}$ is trained for model ensemble. In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense feedback. To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process. After all, the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high accuracy in classifying clean and adversarial images.
Our scenario is established on the $N$-way $K$-shot classification paradigm. Attacks are launched under the gray-box setting [32] in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage. At the $i$-th stage, where $i=1,2,\cdots,T$, a new attack $A_{i}(\cdot)$. is detected and the defender receives a set of feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
@@
In the dynamic environment, various new attacks will emerge on the internet as time goes on. Therefore, the defense mechanism must be capable of adapting to a range of new attacks across different time stages while also retaining knowledge of previous ones.
@@
Over time, the ongoing influx of attacks results in accumulating defense feedback, potentially leading to memory constraints. In practical terms, the defender may not have sufficient memory space to accommodate this accumulation. Consequently, storing all received defense feedback for adaptation is impractical and should be avoided.
@@
The previous defense strategies have led to a sacrifice in the classification accuracy of clean images, resulting in performance degradation. This decline could potentially hurt crucial real-world business operations. As a result, a defender must prioritize maintaining performance on both clean and adversarial images.
@@ 4.1 Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attack using defense feedback $\mathcal{A}_{\mathrm{train}}^{i}$. Inspired by continual learning (CL) [27], we regard classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion.
Following [22], we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage. The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$ and $N_{i}=N\times(i+1)$ is number of classes in stage $i$ after expansion. Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^1] We omit the networks’ parameter in formulas:
@@ 4.2 Embedding Reservation for Few-Shot Adaptation
First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as "virtual classes" [43], where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
@@ 4.3 Prototype Augmentation for Memory Efficiency
In response to Principle 3, we use prototype augmentation to efficiently manage memory.
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [50]. Many CL methods store a fraction of old data to jointly train the model with current data to address this issue [27]. However, preserving old feedback may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation [51] to maintain the decision boundary of previous stages, without saving any feedback. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
@@
4.4 Model Ensemble for Clean and Adversarial Image Classification
The overall algorithm of CAD framework is presented in Algorithm 1.